### v1

# Disclosure of AI-Generated News Increases Engagement but Does Not Reduce Aversion, Despite Positive Quality Ratings

Abstract
The advancement of artificial intelligence (AI) has led to its application in many areas, including journalism. One key issue is the public’s perception of AI-generated content. This preregistered study investigates (i) the perceived quality of AI-assisted and AI-generated versus human-generated news articles, (ii) whether disclosure of AI’s involvement in generating these news articles influences engagement with them, and (iii) whether such awareness affects the willingness to read AI-generated articles in the future. We employed a between-subjects survey experiment with 599 participants from the German-speaking part of Switzerland, who evaluated the credibility, readability, and expertise of news articles. These articles were either written by journalists (control group), rewritten by AI (AI-assisted group), or entirely generated by AI (AI-generated group). Our results indicate that all news articles, regardless of whether they were written by journalists or AI, were perceived to be of equal quality. When participants in the treatment groups were subsequently made aware of AI’s involvement in generating the articles, they expressed a higher willingness to engage with (i.e., continue reading) the articles than participants in the control group. However, they were not more willing to read AI-generated news in the future. These results suggest that aversion to AI usage in news media is not primarily rooted in a perceived lack of quality, and that by disclosing using AI, journalists could attract more immediate engagement with their content, at least in the short term.
1 Introduction
The landscape of journalism is evolving significantly with the advancement of artificial intelligence (AI). Traditionally, journalists have used computers primarily for research and data analysis, but this role is rapidly expanding. Modern AI, particularly generative AI, now matches or even surpasses human capabilities in tasks like text composition. This has led to the emergence of “automated” or “robot” journalism (Carlson, 2018; Clerwall, 2014; Graefe et al., 2018), a key aspect of the broader computational journalism trend that underscores the increasing role of computation and data in news production (Anderson, 2013; Cohen, Hamilton and Turner, 2011; Lewis and Westlund, 2015).
AI applications, especially in natural language generation, have become integral in practical journalism. These applications include news writing for media organizations such as the Associated Press and Forbes, summarizing scientific data via platforms like the Open Research Knowledge Graph, and narrative writing using tools like ChatGPT-4. These advancements enable AI to produce content that is virtually indistinguishable from human writing, demonstrating AI’s potential in various writing-intensive domains.
The integration of AI-generated content in journalism presents several implications. On one hand, it offers opportunities for faster, multilingual, and expanded content production with potentially fewer errors, which could enhance news quality and help combat misinformation (Graefe et al., 2018). AI can also handle routine reporting tasks, allowing journalists to focus on investigative stories. On the other hand, however, there are concerns about job losses in the journalism sector and doubts regarding algorithms’ ability to fulfill the “watchdog” role traditionally associated with journalists (Strömbäck, 2005; Latar, 2015).
A crucial issue is the public’s perception of AI-assisted journalism. Understanding attitudes towards AI-generated journalism is essential as it may directly impact the public’s political knowledge. Indeed, if the public distrusts AI-generated content and chooses not to engage with it, they may miss out on important political information, leading to a less informed citizenry. A recent study revealed that only 29% of Swiss respondents would read fully AI-generated news, 55% would read news if AI assisted in its creation, and 84% would prefer news created without AI involvement (Vogler et al., 2023). This highlights skepticism towards AI-generated news amongst the Swiss population, consistent with surveys from other countries (Fletcher and Nielsen, 2024). Such aversion could be problematic because it could undermine the credibility and effectiveness of news organizations. Therefore, transparent and responsible use of AI in journalism is essential to build trust and ensure that the public remains informed and engaged with credible news sources.
In this preregistered study, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future. Our results indicate that all kinds of news articles, regardless of whether they were written by journalists or (assisted by) AI, were perceived to be of similar quality in terms of credibility, readability, and expertise. When participants in the treatment groups were subsequently informed about AI’s role in generating the articles, they expressed a higher willingness to read the full news article than participants in the control group, who saw human-written articles. This increased engagement, however, did not translate to an increased openness towards reading AI-generated news in the future.
These findings imply that aversion towards AI usage in the news media is not primarily linked to concerns that AI-generated content is of lower quality. Moreover, the results suggest that transparency in the use of AI by journalists could enhance immediate reader engagement. Future work should delve deeper into understanding the underlying mechanisms driving both results, exploring factors such as the novelty of AI involvement, trust in AI, and the psychological effects of transparency. Additionally, research should investigate long-term impacts on reader behavior and trust to develop strategies for effectively integrating AI in journalism, ensuring that it complements rather than undermines the journalistic process.
2 Previous Research
Several strands of literature have explored the perception of human versus AI-generated news. The first strand involves research on public opinions about human versus computer-generated news, relying on respondents’ prior experiences or perceptions (e.g., Vogler et al., 2023; Chuan, Tsai and Cho, 2019). A recent survey conducted across six countries (Argentina, Denmark, France, Japan, the UK, and the USA) revealed that public opinion on the use of generative AI in journalism is mixed (Fletcher and Nielsen, 2024). While people expect AI-produced news to be cheaper and more up-to-date, they also anticipate it to be less trustworthy and transparent. Most respondents are wary of AI-generated news, especially on hard topics like politics and international affairs, but show more acceptance for soft news like fashion and sports. A large majority favors disclosure when AI is used, but opinions differ on which specific uses should be labeled. Uncertainty remains, with many respondents unsure or neutral about AI’s role in journalism. A limitation of these kinds of opinion-based studies is that they could reflect respondents’ preconceptions or lack of familiarity with AI-generated texts, which may not align with actual behaviors when confronted with AI news.
The second strand consists of surveys where respondents directly evaluate news stories labeled as either “human-written” or “computer-written.” For instance, Haim and Graefe (2017) found that respondents rated human-written news higher in readability but lower in credibility compared to AI-generated news. However, when participants were only given a single article without information on its origin, the differences in their evaluations were minimal, indicating that the perception of the author (human or AI) plays as significant a role in the assessment as the content itself.
The third strand focuses on experiments where participants are presented with identical articles but with different stated authors, to measure the effect of the source. For example, Van der Kaa and Krahmer (2014) found that while news consumers rated the credibility of both human and computer-generated articles equally, journalists perceived human-labeled articles as significantly more trustworthy. This effect varied with the topic, as sports articles were rated as less trustworthy than financial ones. Jung et al. (2017) used a $2\times 22 × 2$ design to explore perceptions of article quality based on authorship labels, finding that both journalists and the public rated AI-written articles higher when they were correctly labeled as such. Another experiment by Tandoc Jr, Yao and Wu (2020) varied the attributed author of a human-written text, finding no significant difference in perceived credibility across different labels. Additionally, Altay and Gilardi (2023) explored how labeling news headlines as “AI-generated” affects accuracy ratings and sharing intentions, finding that such labels decreased perceived accuracy and sharing intent, even when the text was human-written.
The fourth strand investigates the effect of the message itself by providing texts written by different authors without stating the source, thereby reducing source-based biases. Clerwall (2014) found that computer-written articles were perceived as more credible, whereas human-written articles scored higher in readability. This approach highlights the challenge of distinguishing between human and AI-generated content. Kim et al. (2020) extended this analysis across multiple mediums (text, audio, video), finding significant differences favoring human-generated content only in audio and video formats. In a follow-up study, Kim, Xu and Merrill Jr (2022) discovered that human newscasters were perceived as more credible than AI newscasters.
Some studies merge the third and fourth strands by measuring both source and message effects. For instance, Graefe et al. (2018) examined the effects of declared source, actual article content, and topic, finding that while declared source effects generally favored human-written articles, computer-written articles scored higher in credibility and expertise in fact-based topics. Wölker and Powell (2021) explored European readers’ perceptions of automated journalism, finding that automated and human content were perceived as equally credible.
Finally, a meta-analysis incorporating results from studies across all four strands found no significant difference in credibility between human-written and AI-generated news, with human-written texts only having a slight advantage in perceived quality and a notable lead in readability (Graefe and Bohlken, 2020). However, all studies in this analysis were conducted before 2020, and the rapid advancements in generative AI may have altered these dynamics.
Overall, existing research suggests that while AI-generated news can be perceived as being of equal quality to human-generated news, knowing that an article was created by AI can negatively influence readers’ perceptions. This implies that preconceived notions may lead to distorted preferences for human-generated news. However, previous studies have not separated quality evaluations from the willingness to read AI-generated news, limiting our understanding of public attitudes in this area. Moreover, recent innovations in AI text generation may have made some of the existing findings outdated, underscoring the need for additional research.
3 Research Questions and Hypotheses
Our study addresses the following research questions:
RQ1:
How do respondents rate the quality (expertise, readability, and credibility) of AI-assisted and AI-generated news articles, compared to human-generated news articles?
RQ2:
What impact does disclosing AI’s role in generating an article have on respondents’ willingness to read AI-assisted or AI-generated news after they have rated the article for quality?
By answering these two research questions, we extend the literature in important ways. First, we combine a detailed measurement of “quality” with three realistic levels of AI-involvement in news generation, reflecting current capabilities and practices. Second, we make sure that respondents’ views on AI are not asked in the abstract, which may reflect inaccurate preconceptions, but are rooted in specific examples we provide. Third, we separate quality evaluations from willingness to read AI-generated news by disclosing the degree of AI involvement after participants evaluated the articles’ quality, which ensures that awareness of AI’s role does not affect quality ratings.
The first set of hypotheses focus on quality assessments of news articles with different degrees of AI involvement. Given that a meta-analysis of experimental studies found that human-written news was generally rated higher for quality (Graefe and Bohlken, 2020), our pre-registered hypotheses state that news articles generated with AI involvement are rated lower for all dimensions of quality. However, we acknowledge that recent advances in generative AI may challenge these expectations.
H1:
Respondents in the AI-assisted group rate the article’s expertise lower than respondents in the control group.
H2:
Respondents in the AI-assisted group rate the article’s readability lower than respondents in the control group.
H3:
Respondents in the AI-assisted group rate the article’s credibility lower than respondents in the control group.
H4:
Respondents in the AI-generated group rate the article’s expertise lower than respondents in the control group.
H5:
Respondents in the AI-generated group rate the article’s readability lower than respondents in the control group.
H6:
Respondents in the AI-generated group rate the article’s credibility lower than respondents in the control group.
Our second set of hypotheses is designed to explore people’s attitudes once they learn the true role of AI in the article they have read, after rating them for quality.
Graefe et al. (2018, 604) found that “news consumers get more pleasure out of reading human-written as opposed to computer-written content”. Hence, we expect that respondents in the AI-assisted and AI-generated groups will show a lower willingness to continue reading their articles after being informed of the true generation process.
H7:
Respondents in the AI-assisted group express a lower willingness to continue reading the article than respondents in the control group.
H8:
Respondents in the AI-generated group express a higher willingness to continue reading the article than respondents in the control group.
Our final hypotheses address attitudes towards AI usage in news articles in general, that is, not regarding the particular articles they have read and rated. Respondents will be asked to what degree they would be willing to read AI-generated news in the future. This approach is more in line with opinion-based studies asking abstract questions. However, in our study we ask this question after respondents have engaged with examples of the kinds of articles we are asking about. There is evidence that greater technological competence, AI familiarity, and AI knowledge increases trust in AI (Novozhilova et al., 2024), which suggests that engagement with AI articles may decrease skepticism towards them. Therefore, we expect that respondents in the AI-assisted and AI-generated display express willingness to read AI-generated news articles in the future.
H9:
Respondents in the AI-assisted group express a higher willingness to read AI-generated news in the future than respondents in the control group.
H10:
Respondents in the AI-generated group express a higher willingness to read AI-generated news in the future than respondents in the control group.
4 Experimental Design
To explore the public’s perceptions of AI-generated and AI-assisted news articles compared to traditional human-generated articles, we conducted an online between-subjects survey experiment. We recruited 599 participants from the age of 18 from the German-speaking part of Switzerland through the survey company Bilendi, using quotas based on age and gender to ensure balance across these dimensions. Sample size was determined by a power analysis, assuming an effect size of Cohen’s d = 0.3 for the primary outcomes and d = 0.15 for the secondary outcomes.
First, participants completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation. Participants were then randomly assigned to one of three conditions: a control group that read excerpts from human-generated articles (“human-generated” group), a first treatment group that read excerpts that were rewritten with the help of ChatGPT (“AI-assisted” group), and another treatment group that read excerpts that were entirely generated by ChatGPT (“AI-generated” group).
To ensure consistency and create a set of texts that is comparable across the three groups, all articles are derived from the human-generated articles that are shown to the control group. We used the following procedure to generate the articles. For AI-assisted articles, we copy-pasted the original article into ChatGPT and asked it to rewrite the article without losing any information. This resulted in articles for the AI-assisted group that are similar to the originals, but are rewritten by the AI system. For AI-generated articles, we only provided ChatGPT with the title and lead of the original article. We then asked ChatGPT to generate a short article in the same style as the source of the human-generated articles.
Each participant read two excerpts, randomly drawn from a pool of ten articles on Swiss politics, with each excerpt having a hard cutoff at 150 words to maintain consistency, avoid bias due to article length, and simulate a paywall for the purposes of our third outcome (willingness to “continue reading” the article). The random selection of two articles from a pool of ten texts for each group minimizes the risk that the outcomes depend on specific topics. Moreover, for each article, we asked respondents to answer a simple question on the article’s content. This allows us to check that respondents have read the text in sufficient detail.
The first outcome measures how participants evaluate the quality of the articles. Following Sundar (1999), we ask respondents to rate three dimensions of quality: expertise, readability, and credibility. Each dimension is based on specific items, expressed as adjectives (Haim and Graefe, 2017): “clear”, ”coherent”, “comprehensive”, “concise”, and “well-written” for expertise; “boring”, “enjoyable”, “interesting”, “lively”, and “pleasing’ for readability; and “biased”, “fair”, and “objective” for credibility. Participants rated the articles on each of these items on a 1-5 Likert-scale, and we then aggregate the scores for each dimension. Following our pre-registered procedure, we then computed a single measure of quality because Cronbach’s alpha was above 0.7. However, results are robust to using separate scores for each dimension.
Following the initial rating, participants were then informed about the true creation process of the articles. After learning how the texts they read were actually created, respondents were asked if they would be willing to continue reading the full articles after reviewing the excerpts again. In addition, all respondents were asked to report if they would be willing to read AI-generated news articles in the future, on a 1-5 Likert-scale.
Following our pre-registration, we use linear regression models to analyze the data, controlling for socio-demographic variables and using the Benjamini-Hochberg procedure to adjust for multiple comparisons.
5 Results
Figure 1 presents the mean ratings for the first outcome, which assesses the perceived quality (credibility, expertise, and readability) of the articles across the three groups: control, AI-assisted, and AI-generated. The mean credibility ratings were 3.39 for the control group, 3.37 for the AI-assisted group, and 3.38 for the AI-generated group. In addition to the minimal variation, the overlapping confidence intervals suggest no significant differences between the groups. For expertise, the mean ratings were 3.56 for the control group, 3.63 for the AI-assisted group, and 3.62 for the AI-generated group, with overlapping confidence intervals again indicating no significant differences. Readability ratings were 3.15 for the control group, 3.14 for the AI-assisted group, and 3.19 for the AI-generated group, with no significant differences observed, as evidenced by the overlapping confidence intervals.
Figure 1: Average ratings (first outcome: quality)
Figure 2: Average ratings (second outcome: willingness to read)
Table 1: OLS regression results (first outcome: quality)
Dependent variable:
Expertise Readability Credibility Expertise Readability Credibility
(H1) (H2) (H3) (H4) (H5) (H6)
AI-assisted 0.049 --0.001 --0.020
(0.066) (0.068) (0.061)
AI-generated 0.073 0.045 --0.010
(0.069) (0.071) (0.062)
Age 0.002 0.007∗∗∗ --0.001 0.002 0.004 0.003
(0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
Gender 0.053 --0.033 --0.056 0.007 0.001 --0.060
(0.068) (0.070) (0.062) (0.071) (0.074) (0.065)
Education 0.001 --0.019 --0.005 0.005 --0.037 0.020
(0.037) (0.039) (0.034) (0.038) (0.040) (0.035)
Political Orientation 0.023 --0.005 --0.016 --0.095∗∗ --0.058 --0.049
(0.035) (0.036) (0.032) (0.037) (0.039) (0.034)
Constant 3.388∗∗∗ 2.892∗∗∗ 3.530∗∗∗ 3.669∗∗∗ 3.168∗∗∗ 3.364∗∗∗
(0.160) (0.165) (0.146) (0.166) (0.173) (0.151)
Observations 358 358 358 345 345 345
R2 0.008 0.029 0.004 0.025 0.017 0.016
Adjusted R2 --0.006 0.015 --0.010 0.011 0.003 0.002
Residual Std. Error 0.623 (df = 352) 0.642 (df = 352) 0.570 (df = 352) 0.633 (df = 339) 0.658 (df = 339) 0.576 (df = 339)
F Statistic 0.575 (df = 5; 352) 2.109 (df = 5; 352) 0.318 (df = 5; 352) 1.750 (df = 5; 339) 1.181 (df = 5; 339) 1.132 (df = 5; 339)
Note: p<<<0.1; ∗∗p<<<0.05; ∗∗∗p<<<0.01. Applying the Benjamini-Hochberg procedure to adjust the p-values for multiple comparisons does not change significance levels.
Table 2: OLS regression results: second outcome (willingness to read)
Dependent variable:
Continue Read Future Read Continue Read Future Read
(1) (2) (3) (4)
AI-assisted 0.561∗∗∗ 0.003
(0.121) (0.122)
AI-generated 0.692∗∗∗ 0.171
(0.126) (0.127)
Age 0.004 --0.018∗∗∗ 0.003 --0.017∗∗∗
(0.004) (0.004) (0.004) (0.004)
Gender --0.048 --0.290∗∗ --0.032 --0.252
(0.124) (0.125) (0.131) (0.133)
Education 0.068 0.130 0.007 0.180∗∗
(0.068) (0.069) (0.070) (0.071)
Political Orientation 0.053 --0.009 --0.081 --0.181∗∗∗
(0.064) (0.065) (0.068) (0.069)
Constant 2.124∗∗∗ 3.067∗∗∗ 2.544∗∗∗ 3.283∗∗∗
(0.291) (0.294) (0.303) (0.308)
Observations 354 357 341 344
R2 0.064 0.072 0.089 0.094
Adjusted R2 0.051 0.059 0.075 0.081
Residual Std. Error 1.130 (df = 348) 1.146 (df = 351) 1.153 (df = 335) 1.173 (df = 338)
F Statistic 4.795∗∗∗ (df = 5; 348) 5.486∗∗∗ (df = 5; 351) 6.512∗∗∗ (df = 5; 335) 7.039∗∗∗ (df = 5; 338)
Note: p<<<0.1; ∗∗p<<<0.05; ∗∗∗p<<<0.01
Table 1 provides a more detailed analysis of the first outcome. The table presents regression coefficients for the AI-assisted and AI-generated groups, with controls for age, gender, education, and political orientation. Each column corresponds to a separate regression model aligned with one of the hypotheses. The coefficients for the AI-assisted and AI-generated groups across expertise, readability, and credibility are not statistically significant, reinforcing the findings from Figure 1 that there are no substantial differences in the perceived quality of the articles between the groups. Among the control variables, only age and political orientation yielded statistically significant coefficients in some models; however, the effect sizes for these variables are very small. These results remain unchanged when estimating the models only for the subset of respondents who passed the manipulation check.
These results indicate that there are no significant differences in perceived credibility, expertise, and readability between AI-assisted, AI-generated, and human-generated articles. Consequently, we reject the first set of hypotheses (H1-H6), which expected lower perceptions of these attributes in the AI-assisted and AI-generated groups compared to the control group. In other words, news articles generated either with the assistance of AI or entirely by AI are perceived to match the quality of traditional articles written by journalists.
Figure 2 illustrates the mean ratings for the second set of outcomes, which measure participants’ willingness to continue reading the articles they were exposed to and their willingness to read AI-generated news in the future. The control group, which read article excerpts written by journalists, had a mean rating of 2.48 for willingness to continue reading, while the AI-assisted group had a rating of 3.08, and the AI-generated group 3.18. The confidence intervals for the AI-assisted and AI-generated groups do not overlap with those of the control group, indicating statistically significant increases in willingness to continue reading for these groups. Additionally, no significant difference is observed between the AI-assisted and AI-generated groups. This leads to the rejection of hypotheses H7 and H8, which predicted lower willingness to continue reading in the AI-assisted and AI-generated groups. For willingness to read AI-generated news in the future, the control group had a mean rating of 2.37, the AI-assisted group 2.34, and the AI-generated group 2.46. The overlapping confidence intervals suggest no significant differences between the groups.
Table 2 provides a detailed overview of the regression results for the second set of outcomes. They reinforce the findings from Figure 2, indicating a higher willingness to continue reading in these groups. For future reading willingness, the coefficients for both the AI-assisted and AI-generated groups are not significant, corroborating the findings from Figure 2 that there is no significant difference in future reading preferences between the different groups. We therefore reject of hypotheses H9 and H10, which predicted higher willingness to read AI-generated news in the future. These results suggest that knowledge of AI involvement and exposure to it do not significantly influence future reading preferences.
In summary, the data show no significant differences in perceived expertise, readability, and credibility between articles created with the help of AI, those fully generated by AI, and those written by humans. However, participants in the AI-assisted and AI-generated groups exhibit a significantly higher willingness to continue reading the articles. These findings suggest that while AI involvement in news production does not negatively impact the perceived quality of news articles when readers are unaware of the original article generation process, it positively influences readers’ willingness to continue reading once the process is revealed. However, no significant effect was detected regarding the willingness to read AI-generated news in the future across the different groups. The rejection of the first set of hypotheses aligns with the notion that modern AI technologies produce content of comparable quality to that of human journalists, while the significant findings in the second outcome highlight an unexpected openness to engaging with AI-generated content, but only in the short term.
6 Conclusion
The primary aim of this paper was to examine the perceived quality of AI-assisted and AI-generated news articles compared to human-generated ones, and to assess how subsequent disclosure of AI involvement influences readers’ willingness to engage with these articles and their openness to AI-generated news in the future. The research design involved a preregistered experimental study where participants evaluated the perceived quality of news articles without knowing their source, followed by disclosure of whether the articles were AI-assisted, AI-generated, or human-written, to measure changes in willingness to engage with the content and future openness to AI-generated news.
We found no significant differences in perceived credibility, expertise, and readability between articles created by AI, those assisted by AI, and those written by humans. However, participants showed a significantly higher willingness to continue reading articles when aware of AI involvement, whether assisted or fully generated, compared to human-written articles. This suggests that AI does not reduce perceived article quality and enhances readers’ engagement once AI origin is disclosed. However, no significant effect was observed regarding future willingness to read AI-generated news, indicating that AI’s role may positively influence immediate engagement, possibly through a novelty effect, but not long-term preferences.
The results regarding our second set of outcomes are particularly noteworthy. We found significant and positive effects on respondents’ willingness to continue reading the articles after the true source was revealed, applicable to both treatment groups (AI-assisted and AI-generated). This suggests that readers may not be opposed to engaging with AI-generated or AI-assisted texts and may even be curious to learn more about AI. However, it is important to consider that the increased interest in reading the full article may not be a lasting effect changing preferences for AI-generated news. Participants might have become curious about AI involvement after learning the true authorship, but they have not become more (nor less) open towards reading AI news in general. Future research could explore this increased interest using more targeted intervention research designs, which could provide valuable insights for news producers on how to enhance specific target audiences’ interest and increase article engagement.
Regarding respondents’ willingness to read AI-generated news articles in the future, we found no significant differences between the treatment groups. Simply reading two AI-assisted or AI-generated texts does not appear to alleviate concerns about AI involvement in news production, as overall willingness scores remained below 2.5 across all groups. These findings may indicate a gradual familiarization process with AI among readers. Over time and with further technological advancements, AI might become more integrated into journalism, particularly in news production, with readers becoming more accustomed to it. Therefore, future research should continue monitoring public attitudes toward AI to detect potential familiarization effects.
Future research could focus on tracking changes in public perception and engagement with AI-generated news over time through longitudinal studies, which would provide insights into whether initial curiosity leads to long-term acceptance or trust. Exploring the role of transparency in disclosing AI involvement would also be important, as different types of transparency might influence how readers perceive and trust AI-generated content. Intervention studies could test different strategies to enhance reader engagement to determine what factors increase willingness to engage with AI-produced news. Expanding this research to include diverse cultural and linguistic contexts would be important, to assess whether these findings are generalizable across different populations. Additionally, studying the impact of AI-generated news on public knowledge could reveal how well readers retain and understand information from AI versus human-written articles.
Acknowledgements
This project received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement nr. 883121). We thank Fabio Melliger for excellent research assistance.

### v2

@@
 Abstract
-The advancement of artificial intelligence (AI) has led to its application in many areas, including journalism.
-One key issue is the public’s perception of AI-generated content.
+The advancement of artificial intelligence (AI) has led to its application in many areas, including news media.
+The integration of AI in journalism presents both opportunities and risks for democracy, making it crucial to understand public reception of and engagement with AI-generated news, as it may directly influence political knowledge and trust.
 This preregistered study investigates (i) the perceived quality of AI-assisted and AI-generated versus human-generated news articles, (ii) whether disclosure of AI’s involvement in generating these news articles influences engagement with them, and (iii) whether such awareness affects the willingness to read AI-generated articles in the future.
@@
 1 Introduction
-The landscape of journalism is evolving significantly with the advancement of artificial intelligence (AI).
-Traditionally, journalists have used computers primarily for research and data analysis, but this role is rapidly expanding.
-Modern AI, particularly generative AI, now matches or even surpasses human capabilities in tasks like text composition.
-This has led to the emergence of “automated” or “robot” journalism (Carlson, 2018; Clerwall, 2014; Graefe et al., 2018), a key aspect of the broader computational journalism trend that underscores the increasing role of computation and data in news production (Anderson, 2013; Cohen, Hamilton and Turner, 2011; Lewis and Westlund, 2015).
-AI applications, especially in natural language generation, have become integral in practical journalism.
-These applications include news writing for media organizations such as the Associated Press and Forbes, summarizing scientific data via platforms like the Open Research Knowledge Graph, and narrative writing using tools like ChatGPT-4.
-These advancements enable AI to produce content that is virtually indistinguishable from human writing, demonstrating AI’s potential in various writing-intensive domains.
-The integration of AI-generated content in journalism presents several implications.
-On one hand, it offers opportunities for faster, multilingual, and expanded content production with potentially fewer errors, which could enhance news quality and help combat misinformation (Graefe et al., 2018).
-AI can also handle routine reporting tasks, allowing journalists to focus on investigative stories.
-On the other hand, however, there are concerns about job losses in the journalism sector and doubts regarding algorithms’ ability to fulfill the “watchdog” role traditionally associated with journalists (Strömbäck, 2005; Latar, 2015).
-A crucial issue is the public’s perception of AI-assisted journalism.
-Understanding attitudes towards AI-generated journalism is essential as it may directly impact the public’s political knowledge.
+The news media landscape is evolving significantly with the advancement of artificial intelligence (AI).
+AI applications, especially in natural language generation, have become integral to the production of news (Cools and Diakopoulos, 2024; Simon, 2024b, a).
+The integration of AI in the news presents both opportunities and risks for democracy (Arguedas and Simon, 2024).
+On the one hand, AI has the potential to make news more accessible and tailored to diverse audiences, supporting informed public debate.
+On the other hand, however, challenges arise around transparency, bias, and trust, as AI models often reflect the dominant worldviews of their creators, potentially narrowing the range of perspectives in public discourse.
+A specific issue is the public’s perception of AI-produced news.
+Understanding attitudes towards AI-generated news is essential as it may directly impact the public’s political knowledge and trust.
 Indeed, if the public distrusts AI-generated content and chooses not to engage with it, they may miss out on important political information, leading to a less informed citizenry.
@@
 Therefore, transparent and responsible use of AI in journalism is essential to build trust and ensure that the public remains informed and engaged with credible news sources.
-In this preregistered study, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future.
+In this preregistered study (see the Appendix for an anonymized version of the pre-registration), conducted with Swiss participants, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future.
 Our results indicate that all kinds of news articles, regardless of whether they were written by journalists or (assisted by) AI, were perceived to be of similar quality in terms of credibility, readability, and expertise.
@@
 This increased engagement, however, did not translate to an increased openness towards reading AI-generated news in the future.
-These findings imply that aversion towards AI usage in the news media is not primarily linked to concerns that AI-generated content is of lower quality.
+These findings imply that aversion towards AI usage in the news media is not primarily linked to concerns that AI-generated content is perceived to be of lower quality.
 Moreover, the results suggest that transparency in the use of AI by journalists could enhance immediate reader engagement.
@@
 2 Previous Research
-Several strands of literature have explored the perception of human versus AI-generated news.
-The first strand involves research on public opinions about human versus computer-generated news, relying on respondents’ prior experiences or perceptions (e.g., Vogler et al., 2023; Chuan, Tsai and Cho, 2019).
-A recent survey conducted across six countries (Argentina, Denmark, France, Japan, the UK, and the USA) revealed that public opinion on the use of generative AI in journalism is mixed (Fletcher and Nielsen, 2024).
-While people expect AI-produced news to be cheaper and more up-to-date, they also anticipate it to be less trustworthy and transparent.
-Most respondents are wary of AI-generated news, especially on hard topics like politics and international affairs, but show more acceptance for soft news like fashion and sports.
-A large majority favors disclosure when AI is used, but opinions differ on which specific uses should be labeled.
-Uncertainty remains, with many respondents unsure or neutral about AI’s role in journalism.
-A limitation of these kinds of opinion-based studies is that they could reflect respondents’ preconceptions or lack of familiarity with AI-generated texts, which may not align with actual behaviors when confronted with AI news.
-The second strand consists of surveys where respondents directly evaluate news stories labeled as either “human-written” or “computer-written.” For instance, Haim and Graefe (2017) found that respondents rated human-written news higher in readability but lower in credibility compared to AI-generated news.
-However, when participants were only given a single article without information on its origin, the differences in their evaluations were minimal, indicating that the perception of the author (human or AI) plays as significant a role in the assessment as the content itself.
-The third strand focuses on experiments where participants are presented with identical articles but with different stated authors, to measure the effect of the source.
-For example, Van der Kaa and Krahmer (2014) found that while news consumers rated the credibility of both human and computer-generated articles equally, journalists perceived human-labeled articles as significantly more trustworthy.
-This effect varied with the topic, as sports articles were rated as less trustworthy than financial ones.
-Jung et al.
-(2017) used a $2\times 22 × 2$ design to explore perceptions of article quality based on authorship labels, finding that both journalists and the public rated AI-written articles higher when they were correctly labeled as such.
-Another experiment by Tandoc Jr, Yao and Wu (2020) varied the attributed author of a human-written text, finding no significant difference in perceived credibility across different labels.
-Additionally, Altay and Gilardi (2023) explored how labeling news headlines as “AI-generated” affects accuracy ratings and sharing intentions, finding that such labels decreased perceived accuracy and sharing intent, even when the text was human-written.
-The fourth strand investigates the effect of the message itself by providing texts written by different authors without stating the source, thereby reducing source-based biases.
-Clerwall (2014) found that computer-written articles were perceived as more credible, whereas human-written articles scored higher in readability.
-This approach highlights the challenge of distinguishing between human and AI-generated content.
-Kim et al.
-(2020) extended this analysis across multiple mediums (text, audio, video), finding significant differences favoring human-generated content only in audio and video formats.
-In a follow-up study, Kim, Xu and Merrill Jr (2022) discovered that human newscasters were perceived as more credible than AI newscasters.
-Some studies merge the third and fourth strands by measuring both source and message effects.
-For instance, Graefe et al.
-(2018) examined the effects of declared source, actual article content, and topic, finding that while declared source effects generally favored human-written articles, computer-written articles scored higher in credibility and expertise in fact-based topics.
-Wölker and Powell (2021) explored European readers’ perceptions of automated journalism, finding that automated and human content were perceived as equally credible.
-Finally, a meta-analysis incorporating results from studies across all four strands found no significant difference in credibility between human-written and AI-generated news, with human-written texts only having a slight advantage in perceived quality and a notable lead in readability (Graefe and Bohlken, 2020).
-However, all studies in this analysis were conducted before 2020, and the rapid advancements in generative AI may have altered these dynamics.
-Overall, existing research suggests that while AI-generated news can be perceived as being of equal quality to human-generated news, knowing that an article was created by AI can negatively influence readers’ perceptions.
+AI is increasingly integrated into news production workflows, handling tasks such as data analysis, trend identification, content aggregation, and personalization (Cools and Diakopoulos, 2024).
+Structured data-based formats like sports and financial reporting are particularly amenable to AI-generated content due to their formulaic nature.
+AI can assist with data-heavy tasks, enabling journalists to focus on responsibilities that demand human judgment, creativity, and critical thinking, such as investigative reporting and nuanced writing (Clerwall, 2014; Haim and Graefe, 2017).
+These applications suggest potential for further AI expansion into other news formats, which may amplify the news industry’s dependency on technological platforms (Simon, 2024b).
+Research on public perceptions of news authored by AI reveals a mixed picture.
+While earlier studies demonstrated preferences for human-written news (Graefe et al., 2018; Graefe and Bohlken, 2020), more recent work found few notable differences in perceived quality, expertise, credibility, and readability between human-written and AI-generated content (Wölker and Powell, 2021; Kim et al., 2020; Kim, Xu and Merrill Jr, 2022).
+However, AI-produced content tend to be distrusted, especially for politically or emotionally sensitive subjects (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020).
+People tend to view AI-produced news as less trustworthy and transparent, and less worth paying for over human-written news (Fletcher and Nielsen, 2024).
+Most support labeling AI involvement, especially for core tasks like writing, with more acceptance in soft news topics than in hard news.
+However, public perceptions of AI-generated news often depend on the specific ways AI is implemented and whether or how its use is disclosed (Altay and Gilardi, 2024).
+AI can be viewed as more objective and less biased than human journalists (Wölker and Powell, 2021; Wang and Ophir, 2024).
+However, perceptions of algorithmic neutrality can be misleading, as AI systems trained on biased data may inadvertently perpetuate or amplify existing biases.
+While transparency is generally seen as beneficial for accountability and audience trust (Wang and Ophir, 2024), simply disclosing AI’s involvement may be insufficient or even counterproductive (Wang and Huang, 2024; Wang and Ophir, 2024).
+People tend to be skeptical of headlines labeled as AI-generated, regardless of whether they are genuinely AI-produced or actually created by humans, partly because they tend to assume an extreme, possibly unrealistic degree of AI involvement (Altay and Gilardi, 2024).
+Without careful management, transparency can lead to negative biases against AI, reducing credibility, particularly among technology-skeptical audiences (Wang and Huang, 2024; Wang and Ophir, 2024).
+AI integration in journalism should be conducted responsibly (Cools and Diakopoulos, 2024).
+Challenges include addressing algorithmic bias, preventing AI-driven misinformation or “hallucinations”, and safeguarding journalistic autonomy and core values (Cools and Diakopoulos, 2024).
+News organizations are encouraged to create structures to experiment with, assess, and refine best practices for incorporating AI tools across all stages of news production, from information gathering to editing and distribution (Araujo et al., 2023; Cools and Diakopoulos, 2024).
+A collaborative model where AI supports rather than replaces journalists, in conjunction with more effective frameworks for sharing information on AI across every level of news organizations, is generally considered best practice (Dodds et al., 2024).
+The integration of AI in journalism presents both opportunities and risks for democracy (Arguedas and Simon, 2024).
+AI offers the promise of enhancing accessibility and personalization in news delivery, fostering more inclusive and informed discussions.
+Yet, concerns persist regarding issues like transparency, bias, and trust, as these systems often mirror the perspectives and biases of their developers, potentially limiting diversity in the viewpoints presented to the public.
+In sum, existing research suggests that while AI-generated news can be perceived as being of equal quality to human-generated news, knowing that an article was created by AI can negatively influence readers’ perceptions.
 This implies that preconceived notions may lead to distorted preferences for human-generated news.
 However, previous studies have not separated quality evaluations from the willingness to read AI-generated news, limiting our understanding of public attitudes in this area.
-Moreover, recent innovations in AI text generation may have made some of the existing findings outdated, underscoring the need for additional research.
+Moreover, recent advancements in AI text generation suggest that some pre-ChatGPT findings may no longer fully apply, emphasizing the importance of continuous, additional research to keep pace with these developments.
 3 Research Questions and Hypotheses
@@
 What impact does disclosing AI’s role in generating an article have on respondents’ willingness to read AI-assisted or AI-generated news after they have rated the article for quality?
-By answering these two research questions, we extend the literature in important ways.
+By answering these two research questions, we extend the literature in three important ways.
 First, we combine a detailed measurement of “quality” with three realistic levels of AI-involvement in news generation, reflecting current capabilities and practices.
@@
 Given that a meta-analysis of experimental studies found that human-written news was generally rated higher for quality (Graefe and Bohlken, 2020), our pre-registered hypotheses state that news articles generated with AI involvement are rated lower for all dimensions of quality.
-However, we acknowledge that recent advances in generative AI may challenge these expectations.
+However, we acknowledge that more recent studies may challenge these expectations (Wölker and Powell, 2021).
 H1:
@@
 Our second set of hypotheses is designed to explore people’s attitudes once they learn the true role of AI in the article they have read, after rating them for quality.
-Graefe et al.
-(2018, 604) found that “news consumers get more pleasure out of reading human-written as opposed to computer-written content”.
-Hence, we expect that respondents in the AI-assisted and AI-generated groups will show a lower willingness to continue reading their articles after being informed of the true generation process.
+Given that AI-generated content is often met with distrust, particularly when it involves political topics (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020), we expect that respondents in the AI-assisted and AI-generated groups will show a lower willingness to continue reading their articles after being informed of the true generation process.
 H7:
@@
 4 Experimental Design
-To explore the public’s perceptions of AI-generated and AI-assisted news articles compared to traditional human-generated articles, we conducted an online between-subjects survey experiment.
+To explore the public’s perceptions of AI-generated and AI-assisted news articles compared to traditional human-generated articles, we conducted an online between-subjects survey experiment (the preregistration containing full study details can be found in the Appendix).
 We recruited 599 participants from the age of 18 from the German-speaking part of Switzerland through the survey company Bilendi, using quotas based on age and gender to ensure balance across these dimensions.
+The study was conducted in May 2024.
 Sample size was determined by a power analysis, assuming an effect size of Cohen’s d = 0.3 for the primary outcomes and d = 0.15 for the secondary outcomes.
-First, participants completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation.
+After informed consent, participants first completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation.
 Participants were then randomly assigned to one of three conditions: a control group that read excerpts from human-generated articles (“human-generated” group), a first treatment group that read excerpts that were rewritten with the help of ChatGPT (“AI-assisted” group), and another treatment group that read excerpts that were entirely generated by ChatGPT (“AI-generated” group).
@@
 Table 2 provides a detailed overview of the regression results for the second set of outcomes.
-They reinforce the findings from Figure 2, indicating a higher willingness to continue reading in these groups.
+They reinforce the findings from Figure 2, indicating a higher willingness to continue reading in the AI-assisted (p<0.01) and AI-generated (p<0.01) groups.
 For future reading willingness, the coefficients for both the AI-assisted and AI-generated groups are not significant, corroborating the findings from Figure 2 that there is no significant difference in future reading preferences between the different groups.

### v3

@@
-Disclosure of AI-Generated News Increases Engagement but Does Not Reduce Aversion, Despite Positive Quality Ratings
+Willingness to Read AI-Generated News Is Not Driven by Their Perceived Quality
 Abstract
-The advancement of artificial intelligence (AI) has led to its application in many areas, including news media.
-The integration of AI in journalism presents both opportunities and risks for democracy, making it crucial to understand public reception of and engagement with AI-generated news, as it may directly influence political knowledge and trust.
+The advancement of artificial intelligence has led to its application in many areas, including news media, which makes it crucial to understand public reception of AI-generated news.
 This preregistered study investigates (i) the perceived quality of AI-assisted and AI-generated versus human-generated news articles, (ii) whether disclosure of AI’s involvement in generating these news articles influences engagement with them, and (iii) whether such awareness affects the willingness to read AI-generated articles in the future.
-We employed a between-subjects survey experiment with 599 participants from the German-speaking part of Switzerland, who evaluated the credibility, readability, and expertise of news articles.
-These articles were either written by journalists (control group), rewritten by AI (AI-assisted group), or entirely generated by AI (AI-generated group).
-Our results indicate that all news articles, regardless of whether they were written by journalists or AI, were perceived to be of equal quality.
-When participants in the treatment groups were subsequently made aware of AI’s involvement in generating the articles, they expressed a higher willingness to engage with (i.e., continue reading) the articles than participants in the control group.
+We conducted a survey experiment with 599 Swiss participants, who evaluated the credibility, readability, and expertise of news articles either written by journalists (control group), rewritten by AI (AI-assisted group), or entirely written by AI (AI-generated group).
+Our results indicate that all articles were perceived to be of equal quality.
+When participants in the treatment groups were subsequently made aware of AI’s role, they expressed a higher willingness to continue reading the articles than participants in the control group.
 However, they were not more willing to read AI-generated news in the future.
-These results suggest that aversion to AI usage in news media is not primarily rooted in a perceived lack of quality, and that by disclosing using AI, journalists could attract more immediate engagement with their content, at least in the short term.
+These results suggest that aversion to AI usage in news media is not primarily rooted in a perceived lack of quality, and that by disclosing using AI, journalists could induce more short-term engagement.
 1 Introduction
@@
 On the other hand, however, challenges arise around transparency, bias, and trust, as AI models often reflect the dominant worldviews of their creators, potentially narrowing the range of perspectives in public discourse.
+Previous research has found that audiences tend to distrust AI-generated content, particularly in politically sensitive contexts, and often prefer human-authored news, even when AI-generated content is evaluated as equally credible (Waddell, 2018; Toff and Simon, 2024; Jia et al., 2024).
 A specific issue is the public’s perception of AI-produced news.
@@
 Such aversion could be problematic because it could undermine the credibility and effectiveness of news organizations.
-Therefore, transparent and responsible use of AI in journalism is essential to build trust and ensure that the public remains informed and engaged with credible news sources.
-In this preregistered study (see the Appendix for an anonymized version of the pre-registration), conducted with Swiss participants, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future.
+While previous research has often suggested that audience reluctance stems from concerns about AI-generated news being of lower quality (Waddell, 2018), this assumption has not been directly tested.
+In this pre-registered study, conducted with Swiss participants, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future.
 Our results indicate that all kinds of news articles, regardless of whether they were written by journalists or (assisted by) AI, were perceived to be of similar quality in terms of credibility, readability, and expertise.
-When participants in the treatment groups were subsequently informed about AI’s role in generating the articles, they expressed a higher willingness to read the full news article than participants in the control group, who saw human-written articles.
-This increased engagement, however, did not translate to an increased openness towards reading AI-generated news in the future.
-These findings imply that aversion towards AI usage in the news media is not primarily linked to concerns that AI-generated content is perceived to be of lower quality.
-Moreover, the results suggest that transparency in the use of AI by journalists could enhance immediate reader engagement.
-Future work should delve deeper into understanding the underlying mechanisms driving both results, exploring factors such as the novelty of AI involvement, trust in AI, and the psychological effects of transparency.
-Additionally, research should investigate long-term impacts on reader behavior and trust to develop strategies for effectively integrating AI in journalism, ensuring that it complements rather than undermines the journalistic process.
-2 Previous Research
-AI is increasingly integrated into news production workflows, handling tasks such as data analysis, trend identification, content aggregation, and personalization (Cools and Diakopoulos, 2024).
+This challenges the assumption that skepticism towards AI-generated journalism is driven by concerns about content quality.
+Instead, our findings align with prior work suggesting that audience resistance is rooted in deeper trust and transparency-related concerns (Barrolleta and Sandoval-Martín, 2024; Bien-Aimé et al., 2025; Jia et al., 2024).
+Moreover, while participants who learned about AI’s involvement showed a higher willingness to continue reading the specific article they were presented with—suggesting an initial engagement effect—this did not translate into a greater willingness to consume AI-generated news in the future.
+This reinforces findings that transparency about AI use can enhance immediate engagement, possibly due to novelty or curiosity, but does not necessarily increase long-term acceptance (Toff and Simon, 2024).
+These results indicate that public aversion to AI-generated news may not be a matter of perceived content quality but rather a broader issue of trust and perception, highlighting the role of audience heuristics in shaping attitudes towards AI in journalism.
+These findings have important implications for both journalism practice and future research.
+If AI-generated news is perceived to be of equal quality but still met with resistance, efforts to increase public acceptance may need to focus on improving trust rather than simply enhancing AI’s capabilities.
+Future work should explore the mechanisms behind this resistance, examining how AI transparency, human oversight, and audience familiarity with AI influence perceptions of credibility.
+Additionally, research should investigate long-term trends in engagement and trust to better understand whether exposure to AI-generated news over time leads to greater acceptance or sustained skepticism.
+Understanding these dynamics will be crucial for integrating AI into journalism in a way that complements rather than undermines public trust in news media.
+2 Perceptions and Challenges of AI-Generated News
+AI is increasingly integrated into news production workflows, handling tasks such as data analysis, trend identification, content aggregation, and personalization (Cools and Koliska, 2024).
 Structured data-based formats like sports and financial reporting are particularly amenable to AI-generated content due to their formulaic nature.
 AI can assist with data-heavy tasks, enabling journalists to focus on responsibilities that demand human judgment, creativity, and critical thinking, such as investigative reporting and nuanced writing (Clerwall, 2014; Haim and Graefe, 2017).
-These applications suggest potential for further AI expansion into other news formats, which may amplify the news industry’s dependency on technological platforms (Simon, 2024b).
+However, concerns persist regarding the opacity of AI-driven processes.
+Research on The Washington Post’s use of AI-driven news automation shows that while external transparency efforts exist, internal newsroom transparency remains limited, which can lead to errors and hinder journalists’ ability to critically assess AI-generated content (Cools and Koliska, 2024).
+These applications suggest potential for further AI expansion into other news formats, which may amplify the news industry’s dependency on technological platforms.
 Research on public perceptions of news authored by AI reveals a mixed picture.
 While earlier studies demonstrated preferences for human-written news (Graefe et al., 2018; Graefe and Bohlken, 2020), more recent work found few notable differences in perceived quality, expertise, credibility, and readability between human-written and AI-generated content (Wölker and Powell, 2021; Kim et al., 2020; Kim, Xu and Merrill Jr, 2022).
-However, AI-produced content tend to be distrusted, especially for politically or emotionally sensitive subjects (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020).
-People tend to view AI-produced news as less trustworthy and transparent, and less worth paying for over human-written news (Fletcher and Nielsen, 2024).
-Most support labeling AI involvement, especially for core tasks like writing, with more acceptance in soft news topics than in hard news.
-However, public perceptions of AI-generated news often depend on the specific ways AI is implemented and whether or how its use is disclosed (Altay and Gilardi, 2024).
-AI can be viewed as more objective and less biased than human journalists (Wölker and Powell, 2021; Wang and Ophir, 2024).
-However, perceptions of algorithmic neutrality can be misleading, as AI systems trained on biased data may inadvertently perpetuate or amplify existing biases.
-While transparency is generally seen as beneficial for accountability and audience trust (Wang and Ophir, 2024), simply disclosing AI’s involvement may be insufficient or even counterproductive (Wang and Huang, 2024; Wang and Ophir, 2024).
+A Turing test-based study assessing Spanish-language AI-generated news similarly found that AI-authored articles were perceived as equally credible to those written by professional journalists, though human authors were still seen as producing more engaging narratives (Barrolleta and Sandoval-Martín, 2024).
+Nonetheless, AI-produced content tends to be distrusted, especially for politically or emotionally sensitive subjects (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020).
+Experimental evidence indicates that when news is explicitly labeled as AI-generated, it is often judged as less trustworthy, even when the content itself is evaluated as accurate or fair (Toff and Simon, 2024; Jia et al., 2024).
+This effect is particularly pronounced among audiences who already hold high levels of trust in news, suggesting that those most engaged with journalism are also most resistant to AI involvement (Toff and Simon, 2024).
+Attribution and transparency play a crucial role in shaping audience perceptions of AI-generated news.
+Research on how readers interpret AI involvement in news articles shows that many assume an exaggerated degree of automation, even when AI plays only a minor role (Bien-Aimé et al., 2025).
+Furthermore, the way AI authorship is framed in bylines significantly influences credibility assessments; studies have found that the more AI appears to contribute to an article, the lower the perceived message and source credibility (Jia et al., 2024).
+However, transparency cues such as source disclosure can mitigate some of these negative effects, especially when AI-generated content is accompanied by explicit references to human oversight or lists of data sources used in its creation (Toff and Simon, 2024).
+The issue of bias in AI-generated news remains a persistent concern.
+While some studies suggest AI is perceived as more objective than human journalists (Wölker and Powell, 2021; Wang and Ophir, 2024), this perception may be misleading.
+AI systems trained on biased data can inadvertently perpetuate and amplify existing biases.
+Experiments have found that audiences tend to evaluate AI-generated news as less biased, but this effect diminishes when AI involvement is explicitly disclosed, suggesting that perceived neutrality is contingent on a lack of transparency rather than an inherent feature of AI-generated content (Waddell, 2019, 2018).
+While transparency is generally seen as beneficial for accountability and audience trust, simply disclosing AI’s involvement may be insufficient or even counterproductive (Wang and Huang, 2024; Wang and Ophir, 2024).
 People tend to be skeptical of headlines labeled as AI-generated, regardless of whether they are genuinely AI-produced or actually created by humans, partly because they tend to assume an extreme, possibly unrealistic degree of AI involvement (Altay and Gilardi, 2024).
-Without careful management, transparency can lead to negative biases against AI, reducing credibility, particularly among technology-skeptical audiences (Wang and Huang, 2024; Wang and Ophir, 2024).
-AI integration in journalism should be conducted responsibly (Cools and Diakopoulos, 2024).
-Challenges include addressing algorithmic bias, preventing AI-driven misinformation or “hallucinations”, and safeguarding journalistic autonomy and core values (Cools and Diakopoulos, 2024).
+Research has also shown that AI-generated content is often seen as more credible when audiences are unaware of its origin, but once labeled as AI-generated, perceptions shift negatively (Jia et al., 2024; Waddell, 2018).
+This phenomenon suggests that audience heuristics about AI play a crucial role in shaping credibility assessments.
+AI integration in journalism should be conducted responsibly (Cools and Koliska, 2024).
+Challenges include addressing algorithmic bias, preventing AI-driven misinformation or “hallucinations,” and safeguarding journalistic autonomy and core values.
+Recent findings highlight the need for clearer internal guidelines on AI use within newsrooms, as many organizations still lack standardized attribution policies (Bien-Aimé et al., 2025).
 News organizations are encouraged to create structures to experiment with, assess, and refine best practices for incorporating AI tools across all stages of news production, from information gathering to editing and distribution (Araujo et al., 2023; Cools and Diakopoulos, 2024).
@@
 3 Research Questions and Hypotheses
-Our study addresses the following research questions:
+Building on the existing literature and the gaps identified in previous studies, our research aims to systematically examine how AI involvement in news generation affects audience perceptions.
+Specifically, we address the following research questions:
 RQ1:
-How do respondents rate the quality (expertise, readability, and credibility) of AI-assisted and AI-generated news articles, compared to human-generated news articles?
+Prior to disclosing AI’s role, how do respondents evaluate the quality of AI-assisted and AI-generated news articles in terms of expertise, readability, and credibility, compared to human-generated news articles?
 RQ2:
-What impact does disclosing AI’s role in generating an article have on respondents’ willingness to read AI-assisted or AI-generated news after they have rated the article for quality?
-By answering these two research questions, we extend the literature in three important ways.
-First, we combine a detailed measurement of “quality” with three realistic levels of AI-involvement in news generation, reflecting current capabilities and practices.
-Second, we make sure that respondents’ views on AI are not asked in the abstract, which may reflect inaccurate preconceptions, but are rooted in specific examples we provide.
-Third, we separate quality evaluations from willingness to read AI-generated news by disclosing the degree of AI involvement after participants evaluated the articles’ quality, which ensures that awareness of AI’s role does not affect quality ratings.
-The first set of hypotheses focus on quality assessments of news articles with different degrees of AI involvement.
-Given that a meta-analysis of experimental studies found that human-written news was generally rated higher for quality (Graefe and Bohlken, 2020), our pre-registered hypotheses state that news articles generated with AI involvement are rated lower for all dimensions of quality.
-However, we acknowledge that more recent studies may challenge these expectations (Wölker and Powell, 2021).
+How does disclosing AI’s role in generating an article influence respondents’ willingness to engage with AI-assisted or AI-generated news after they have rated the article’s quality?
+These questions allow us to extend prior research in three ways.
+First, we operationalize “quality” as a multidimensional construct, assessing expertise, readability, and credibility across three distinct levels of AI involvement, reflecting real-world journalistic practices.
+Second, instead of relying on abstract survey questions about AI, which may be influenced by preconceptions, we ground participants’ responses in concrete examples of AI-generated and AI-assisted news.
+Third, we separate the evaluation of news article quality from attitudes toward AI-generated news by only revealing AI’s involvement extitafter respondents have rated the articles.
+This design ensures that initial quality assessments are not biased by prior knowledge of AI authorship.
+3.1 Quality Perceptions of AI-Generated News
+The first set of hypotheses investigates whether AI-assisted or AI-generated news is perceived as lower in quality compared to human-written articles.
+Prior studies have found that human-written news is generally rated higher for quality (Graefe and Bohlken, 2020).
+This aligns with the authority heuristic and social presence heuristic, which suggest that content created by humans is typically seen as more credible due to its perceived human expertise and social presence (Sundar, 2008).
+Given these established biases, we expect that articles with AI involvement will receive lower quality ratings, though more recent findings challenge this assumption (Wölker and Powell, 2021; Barrolleta and Sandoval-Martín, 2024).
 H1:
@@
 Respondents in the AI-generated group rate the article’s credibility lower than respondents in the control group.
-Our second set of hypotheses is designed to explore people’s attitudes once they learn the true role of AI in the article they have read, after rating them for quality.
-Given that AI-generated content is often met with distrust, particularly when it involves political topics (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020), we expect that respondents in the AI-assisted and AI-generated groups will show a lower willingness to continue reading their articles after being informed of the true generation process.
+3.2 Impact of AI Disclosure on Engagement
+The second set of hypotheses examines how awareness of AI involvement affects engagement with the article.
+Previous research suggests that AI-generated content is often met with skepticism, particularly when it covers political topics (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020).
+We therefore hypothesize that respondents will be less inclined to continue reading once they learn about AI’s role.
+However, recent findings that disclosure may also spark curiosity or novelty-driven engagement may challenge this premise (Toff and Simon, 2024).
 H7:
@@
 Respondents in the AI-generated group express a higher willingness to continue reading the article than respondents in the control group.
-Our final hypotheses address attitudes towards AI usage in news articles in general, that is, not regarding the particular articles they have read and rated.
-Respondents will be asked to what degree they would be willing to read AI-generated news in the future.
-This approach is more in line with opinion-based studies asking abstract questions.
-However, in our study we ask this question after respondents have engaged with examples of the kinds of articles we are asking about.
-There is evidence that greater technological competence, AI familiarity, and AI knowledge increases trust in AI (Novozhilova et al., 2024), which suggests that engagement with AI articles may decrease skepticism towards them.
-Therefore, we expect that respondents in the AI-assisted and AI-generated display express willingness to read AI-generated news articles in the future.
+3.3 Long-Term Attitudes Toward AI-Generated News
+The final set of hypotheses assesses how engagement with AI-generated content influences respondents’ general willingness to read AI-generated news in the future.
+Unlike the previous hypotheses, which examine reactions to specific articles, these hypotheses explore broader attitudes towards AI-generated journalism.
+While prior surveys indicate high levels of skepticism towards AI-generated news (Vogler et al., 2023), experimental research suggests that direct exposure to AI-generated content may mitigate these concerns, as familiarity with AI often reduces skepticism and increases trust (Novozhilova et al., 2024).
 H9:
@@
 Respondents in the AI-generated group express a higher willingness to read AI-generated news in the future than respondents in the control group.
+By disentangling quality assessments from short-term engagement dynamics and long-term attitudes, our study contributes to a more detailed understanding of how AI-generated journalism is perceived.
 4 Experimental Design
@@
 Sample size was determined by a power analysis, assuming an effect size of Cohen’s d = 0.3 for the primary outcomes and d = 0.15 for the secondary outcomes.
-After informed consent, participants first completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation.
+Figure 1: Survey experiment flow
+Figure 1 depicts the survey experiment flow: After informed consent, participants first completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation.
 Participants were then randomly assigned to one of three conditions: a control group that read excerpts from human-generated articles (“human-generated” group), a first treatment group that read excerpts that were rewritten with the help of ChatGPT (“AI-assisted” group), and another treatment group that read excerpts that were entirely generated by ChatGPT (“AI-generated” group).
-To ensure consistency and create a set of texts that is comparable across the three groups, all articles are derived from the human-generated articles that are shown to the control group.
-We used the following procedure to generate the articles.
+To ensure consistency and create a set of texts that is comparable across the three groups, all articles are derived from actual articles on Swiss politics published on the news website of Switzerland’s public broadcaster (SRF), which ensures homogeneity in style, quality, as well as topic to a certain extent (that is, only Swiss politics, with variation in sub-topics).
+These articles are shown to the control group.
+We used the following procedure to generate articles with AI involvement.
 For AI-assisted articles, we copy-pasted the original article into ChatGPT and asked it to rewrite the article without losing any information.
@@
 We then asked ChatGPT to generate a short article in the same style as the source of the human-generated articles.
-Each participant read two excerpts, randomly drawn from a pool of ten articles on Swiss politics, with each excerpt having a hard cutoff at 150 words to maintain consistency, avoid bias due to article length, and simulate a paywall for the purposes of our third outcome (willingness to “continue reading” the article).
+Each participant read two excerpts, randomly drawn from a pool of ten articles for each category (human-generated, AI-assisted, AI-generated), making a total pool of 30 articles across all experimental conditions.
+Each excerpt had a hard cutoff at 150 words to maintain consistency, avoid bias due to article length, and simulate a paywall for the purposes of our third outcome (willingness to “continue reading” the article).
 The random selection of two articles from a pool of ten texts for each group minimizes the risk that the outcomes depend on specific topics.
@@
 However, results are robust to using separate scores for each dimension.
-Following the initial rating, participants were then informed about the true creation process of the articles.
-After learning how the texts they read were actually created, respondents were asked if they would be willing to continue reading the full articles after reviewing the excerpts again.
+In a second step — following the initial rating — participants were informed about the actual creation process of the articles.
+After learning how the texts they read were created, respondents were asked if they would be willing to continue reading the full articles after reviewing the excerpts again.
 In addition, all respondents were asked to report if they would be willing to read AI-generated news articles in the future, on a 1-5 Likert-scale.
@@
 This suggests that readers may not be opposed to engaging with AI-generated or AI-assisted texts and may even be curious to learn more about AI.
+This finding is in contrast with the concern that readers would disapprove of automated news, leading news organizations to refrain from disclosing that a story was automatically generated (Graefe and Bohlken, 2020), yet mirrors recent research showing that recognizing the benefits of AI in journalism, such as improved objectivity and accuracy, positively influences trust and engagement (Morosoli et al., 2024).
 However, it is important to consider that the increased interest in reading the full article may not be a lasting effect changing preferences for AI-generated news.
 Participants might have become curious about AI involvement after learning the true authorship, but they have not become more (nor less) open towards reading AI news in general.
-Future research could explore this increased interest using more targeted intervention research designs, which could provide valuable insights for news producers on how to enhance specific target audiences’ interest and increase article engagement.
+Future research could explore this increased interest using more targeted intervention research designs.
+This could provide valuable insights for news producers on how to enhance specific target audiences’ interest and increase article engagement, especially now that numerous news organizations are setting up guidelines to govern how they use AI (Becker, Simon and Crum, 2025).
 Regarding respondents’ willingness to read AI-generated news articles in the future, we found no significant differences between the treatment groups.
@@
 Over time and with further technological advancements, AI might become more integrated into journalism, particularly in news production, with readers becoming more accustomed to it.
-Therefore, future research should continue monitoring public attitudes toward AI to detect potential familiarization effects.
+However, our study suggests that audience hesitation is not purely a matter of unfamiliarity but is also tied to deeper normative expectations about journalistic authenticity.
+Beyond its empirical findings, this study contributes to theoretical discussions on media trust, audience behavior, and the evolving role of AI in journalism.
+Notably, our results indicate a decoupling of content quality from audience acceptance, reinforcing the idea that trust in journalism is not solely determined by article quality but also by perceptions of authorship.
+This aligns with the concept of algorithmic aversion, wherein readers are reluctant to accept AI-generated content despite recognizing its quality (Dietvorst, Simmons and Massey, 2015).
+The study extends source credibility theory by demonstrating that while AI-generated articles may score high in expertise, they may still be viewed as less trustworthy due to the absence of human editorial oversight, accountability, and perceived authenticity (Altay and Gilardi, 2024).
+These findings also contribute to ongoing debates about transparency in journalism.
+While transparency is often promoted as a mechanism to build trust, our results suggest that disclosing AI involvement in news production does not necessarily lead to increased long-term acceptance, highlighting the need for more refined models of how AI disclosure interacts with audience heuristics and expectations.
+Future research should investigate how different framing strategies—such as emphasizing human oversight or explaining AI’s editorial role—might mitigate trust concerns.
 Future research could focus on tracking changes in public perception and engagement with AI-generated news over time through longitudinal studies, which would provide insights into whether initial curiosity leads to long-term acceptance or trust.
@@
 Additionally, studying the impact of AI-generated news on public knowledge could reveal how well readers retain and understand information from AI versus human-written articles.
+Finally, our findings emphasize the importance of balancing AI integration in journalism with audience expectations, suggesting that future news production models may benefit from hybrid approaches that maintain a human editorial presence while leveraging AI for efficiency.
+Ethics Approval
+This study was approved by the Ethics Committee of the Faculty of Arts and Social Sciences, University of Zurich (approval number 23.10.14).
 Acknowledgements